im havin trouble with this equation can you solve it and show me wat to do
5p+8(3p+4)=264

Answer by fractalier(6550)

Okay from
5p + 8(3p + 4) = 264
let's expand, collect terms and solve...like this
5p + 24p + 32 = 264
29p + 32 = 264 subtract 32 from both sides
29p = 232 divude by 29 and get
p = 8